[Hallc_running] RC Daily Report

Michael Paolone tue81317 at temple.edu
Sun Mar 3 15:38:35 EST 2019

Hi Everyone,

Around 18:30 last night, beam was restored and delivered to Hall-C at a
maximum of 50uA.  Due to the extended time lost yesterday, beam studies
will be truncated from all day-shift to 8:30 - noon.  We are still on
kinematic-4, and still plan to stay with it through the end of the
experiment.  There are no planned changes or accesses needed


- At 23:00 we had an SHMS dipole interlock/slow-discharge and at 6:00 we
had the same issue with Q3.  In both cases, Steve Lassiter was contacted,
and he was able to fix the issue remotely.  We were able to ramp-up and
recover without additional issues both times. See logbook entries
https://logbooks.jlab.org/entry/3662259 and
- We are having target white alarms every 3 hours or so.  It seems whatever
was done to fix this before ( https://logbooks.jlab.org/entry/3660122), is
no longer fixing the problem.

There is no need for a 4:30 meeting today.  We will have one as usual on

-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20190303/50dc367d/attachment.html>

More information about the Hallc_running mailing list